# Path to Diversity: A Primer on ISAC-izing Commodity Wi-Fi for Practical Deployments

### v1

@@ I-A The ISAC Paradigm and the Wi-Fi Potential
Among the various technologies supporting ISAC, commodity Wi-Fi stands out as a practical and promising platform. This prominence is primarily driven by its ubiquity and low deployment cost—intrinsic strengths that render it an ideal foundation for scalable sensing without requiring dedicated infrastructure [51]. Since Channel State Information (CSI) [18, 17] became explicitly accessible, Wi-Fi has evolved into a versatile sensor supporting diverse device-free applications, ranging from localization [37, 87] and activity recognition [32, 11] to vital sign monitoring [50, 99]. Furthermore, distinct from vision-based modalities, Wi-Fi sensing inherently preserves user privacy and ensures robustness against poor lighting and severe occlusion [60, 90]. These properties establish Wi-Fi as the most accessible entry point for real-world ISAC deployments.
@@ I-D Existing Survey and Tutorial Landscape
Jiang et al. (2018) [31] Application-Centric Focuses on smart home use cases; classifies Wi-Fi sensing tasks into health, interaction, authentication, etc. Application-driven but signal-agnostic; does not examine physical constraints or capabilities.
Xiao et al. (2023) [84] Application-Centric Proposes a dual-axis taxonomy (task type × motion granularity); reviews diverse human sensing scenarios. High-level abstraction of application logic; lacks analysis of signal-layer capabilities or diversity mechanisms.
He et al. (2020) [20] Application-Centric Frames Wi-Fi sensing as “Wi-Fi vision”; categorizes perception-level tasks such as imaging, detection, and pose. Uses vision-based analogies; lacks generalizable signal-layer principles; diversity implicit but not abstracted.
Li et al. (2022) [39] Algorithm-Centric Reviews deep learning architectures across sensing pipeline stages; discusses model performance and learning strategies. Strong focus on DL methods; minimal engagement with signal formation, bandwidth effects, or spatial richness.
Yang et al. (2023) [94] Algorithm-Centric + Benchmarking Builds SenseFi, a benchmarking library for comparing DL models across sensing tasks and datasets. Empirical emphasis on model/dataset comparison; abstracted from signal-level or physical-layer sensing implications.
Chen et al. (2023) [3] Algorithm-Centric Focuses on cross-domain generalization techniques (e.g., transfer learning, domain-invariant features) for CSI-based sensing. Centered on inference transferability; does not address signal-side factors affecting domain shift or generalization.
Ahmad et al. (2024) [1] Algorithm-Centric Surveys deep learning models (CNN, LSTM, Transformers) and training strategies for Wi-Fi-based human sensing; discusses system-level deployment challenges and future DL directions. Strong DL focus; limited discussion of signal-level diversity, CSI formation, or physical-layer modeling assumptions.
Ma et al. (2020) [51] Signal-Level Surveys CSI signal modeling, preprocessing, and traditional sensing algorithms; introduces signal feature taxonomies. Covers signal chain in depth but lacks unified diversity abstraction; diversity dimensions only appear implicitly.
Liu et al. (2020) [49] Broad / Mixed Provides a general overview of wireless sensing modalities (CSI, RSSI, FMCW) and their applications. Broad but shallow; signal-level discussion lacks depth and abstraction; no synthesis of diversity roles.
Hernandez & Bulut (2023) [26] System-Centric Reviews signal processing and edge-deployment challenges for practical Wi-Fi sensing; emphasizes resource constraints. Deployment-focused; diversity viewed through hardware bottlenecks, not as a sensing-enabling mechanism.
@@
I-D1 Application-Centric Surveys
@@
These surveys primarily organize the literature by sensing task types or deployment scenarios, offering taxonomies that are practical for developers or system designers. However, they typically treat the Wi-Fi physical layer as an abstracted sensing interface. Jiang et al. (2018) [31] is among the earliest surveys in this direction, focusing on Wi-Fi-based smart home applications. It organizes use cases into functional categories such as health monitoring, user authentication, and gesture control. He et al. (2020) [20] frames Wi-Fi sensing through the analogy of “Wi-Fi vision,” categorizing tasks such as imaging, detection, and pose estimation in sensing-level terms. Xiao et al. (2023) [84] proposes a dual-axis taxonomy structured by sensing task type and motion granularity, offering a broad mapping between human behaviors and sensing systems. While these works provide valuable perspectives on high-level task types, they offer limited insight into the underlying signal structure or physical-layer enablers. They tend to discuss what can be sensed, while abstracting away how the physical signal diversity constrains or enables these capabilities.
I-D2 Algorithm-Centric and Learning-Oriented Surveys
These surveys focus on machine learning models, training paradigms, and inference frameworks, often treating Channel State Information (CSI) as a given input modality. Li et al. (2022) [39] reviews deep learning architectures across different stages of the Wi-Fi sensing pipeline—from data preprocessing and feature representation to inference models. Yang et al. (2023) [94] develops SenseFi, a benchmarking library that compares DL models across different datasets and sensing tasks. Chen et al. (2023) [3] surveys techniques for cross-domain generalization, such as transfer learning and domain-invariant representations, to improve robustness under environmental or user variation. Most recently, Ahmad et al. (2024) [1] provides a unified overview of deep learning for Wi-Fi sensing, covering model structures, training strategies, deployment challenges, and open research opportunities. While providing rigorous algorithmic taxonomies, these works generally decouple the learning process from the signal formation mechanism. They tend to frame sensing challenges—such as low resolution or poor generalization—as problems to be solved solely through model complexity or data augmentation. Consequently, they overlook how physical-layer diversity fundamentally dictates the information content and the theoretical performance limit of any downstream algorithm.
I-D3 Signal-Level and System-Centric Surveys
A third line of surveys examines signal modeling, CSI extraction, and system implementation challenges. These works engage more directly with the physical-layer properties that underlie Wi-Fi sensing. Ma et al. (2020) [51] systematically reviews CSI signal modeling, feature extraction, and algorithmic processing pipelines for various sensing tasks. Liu et al. (2020) [49] provides a broad overview of RF sensing modalities—including CSI, RSSI, and FMCW—linking each to high-level application types. Hernandez and Bulut (2023) [26] address signal processing and machine learning deployment on edge devices, with a particular emphasis on computational efficiency and system bottlenecks. While engaging with physical-layer details, these works lack a unified theoretical abstraction. They typically treat signal parameters as isolated hardware specifications or implementation hurdles, rather than systematically analyzing how these resources collectively determine the fundamental limits of sensing.
@@ I-E Tutorial Motivation, Contribution, and Structure
  • Sec. VII and Sec. VIII discuss future directions (e.g., Native ISAC, AI foundation models) and conclude the tutorial.
@@ II-A1 Physical Principles and Mathematical Representation of CSI
Channel State Information (CSI) provides a fine-grained characterization of the physical layer of a wireless communication channel. Conceptually, it is a discrete, sampled measurement of the underlying Channel Frequency Response (CFR), which is the continuous function that describes how a signal’s amplitude and phase change across all frequencies. In modern Wi-Fi systems based on Orthogonal Frequency-Division Multiplexing (OFDM), CSI is estimated for each subcarrier from known training sequences, such as the Long Training Field (LTF), embedded in each data packet’s preamble.
@@ II-B1 The Reality of CSI: Hardware-Induced Offsets
Modeling frameworks typically decompose these hardware-induced offsets into four major categories, each originating from specific components of the RF and baseband processing chain [86]. These include Carrier Frequency Offset (CFO), Carrier Phase Offset (CPO), Sampling Frequency Offset (SFO), and Packet Detection Delay (PDD). While all four manifest as phase distortions, they differ fundamentally in their temporal characteristics and impact on sensing.
@@ II-E2 Physical Essence (Mapping Feedback to Spatial Beams)
However, stability comes at the cost of fidelity. Spatially, BFI discards amplitude information ($\bm{\Sigma}$) and absolute phase, rendering it ill-suited for attenuation-based ranging. Temporally, unlike continuous CSI streams, real-world BFI is inherently sparse and bursty. It is generated only during traffic-dictated sounding intervals, leading to irregular sampling rates that challenge continuous tracking—a bottleneck addressed by advanced recovery algorithms in Sec. V-D.
@@ III-C2 Performance Evaluation and Capabilities
Figure 8: Ranging error comparison between ISAC-Fi (Monostatic) and traditional baselines (MUSIC/IFFT). By eliminating synchronization offsets via a shared clock, ISAC-Fi achieves significantly lower and more stable ranging errors [7].
@@ III-D Summary : Towards Synchronized ISAC
This chapter demonstrated how Temporal Diversity is reclaimed in Wi-Fi sensing by resolving the synchronization-interference dilemma. We identified that while bistatic setups suffer from clock asynchrony (losing absolute accuracy), monostatic setups suffer from self-interference (losing visibility). To address these challenges, we have categorized the engineering solutions into two distinct pathways:
  • Active Self-Interference Cancellation: This methodology employs a multi-stage pipeline (combining analog and digital cancellation) to actively subtract the transmission leakage. By recovering the complete channel response, it enables true absolute ranging for both static and dynamic targets. As demonstrated by the ISAC-Fi system [7], this approach offers the highest sensing fidelity but entails significant hardware complexity.
  • Physical Signal Separation: This methodology prevents interference at the source through hardware isolation and self-mixing architectures. It achieves high-precision Phase Coherence by eliminating common-mode clock errors. As exemplified by SiWiS [61], this approach excels in robust motion tracking with lower computational overhead, though its reliance on DC removal typically restricts it to dynamic sensing.
Table IV provides a comparative analysis of these two fundamental approaches. Ultimately, these technologies transform the Wi-Fi physical layer from a simple communication interface into a precision timing instrument, effectively achieving Synchronized ISAC. With the capability to measure range and motion accurately established, the next challenge lies in improving the resolution of these measurements—a task governed by Frequency Diversity, which we explore in the next chapter.
@@ IV-D2 The Reconstruction Challenge
Primary Sensing Output High-Resolution PDP or ToA ToF of the Direct Path High-Resolution ToF-AoA Spectrum
@@ IV-E Summary: Towards High-Resolution ISAC
This chapter charted the evolution of technologies designed to break the physical bandwidth limit of commodity Wi-Fi, a journey aimed at realizing High-Resolution ISAC. We established that achieving this goal fundamentally requires synthesizing a wide effective bandwidth to minimize the range resolution ($\Delta\tau=1/B$). The field has progressed from early attempts at physically stitching adjacent channels to the modern paradigm of discrete, computational reconstruction.
As summarized in Table V, this evolution represents a shift from hardware-centric constraints to algorithmic innovation:
  • Continuous Channel Stitching: Relies on stitching physically contiguous channels. While intuitive, it is severely hindered by the scarcity of contiguous spectrum and the influence of boundary effect [86].
  • Multi-Band Frequency Sweeping: Introduces non-stitching sweeping to resolve range ambiguity. However, its reliance on a sequential multi-band sweep and active protocols limits its applicability to general sensing [69].
  • Discrete Channel Sampling: Combines fast, randomized channel hopping with model-driven deep learning. It synthesizes a 4.7 GHz aperture from sparse discrete samples, successfully unlocking centimeter-level resolution and general sensing capabilities on commodity hardware [45].
With fine-grained resolution established via Frequency Diversity, another remaining challenge is spatial coverage. A single viewpoint, no matter how wide its bandwidth, is vulnerable to blockage and occlusion. This motivates the exploration of Link Diversity in the following chapter, discussing how multi-perspective observations enable robust, distributed ISAC.
@@ V Link Diversity: Realizing Ubiquitous ISAC via Physical Multi-Links and Digital Feedback
Paradigm Core Mechanism Representative Works Key Capability
Coverage & Triangulation Aggregates views to eliminate blind spots; exploits AoA, shadowing, and backscatter for geometric solving. WiSee [53], E-eyes [77], WiDraw [63], LiFS [74], WiDeo [36] Robustness:
No blind spots.
View-Invariant Feature Extraction Combines multi-link Doppler to reconstruct absolute velocity vectors, or extracts BVP for domain independence. IndoTrack [41], WiTraj [83], RIM [13], WiDance [56], Widar3.0 [102], DeepMV [92] Generalization:
View/Location invariance.
3D Structure recovery.
Multi-user separation.
@@ V-B1 Omnidirectional Coverage and Geometric Triangulation
The primary objective is to eliminate inherent blind spots by aggregating views. This philosophy originates from  [97], which utilized distributed links to probabilistically infer user presence. Building on this coverage capability, WiSee [53] leveraged Doppler shifts for whole-home gesture recognition, while E-eyes [77] used fine-grained WiFi signatures from home appliances to identify activities. Beyond detection, researchers aimed for finer geometric solving. WiDraw [63] enables hands-free drawing by exploiting the Angle-of-Arrival (AoA) and hand occlusions to track fine-grained trajectories. In parallel, LiFS [74] targets room-scale localization with low human effort, modeling CSI measurements as power fading equations to solve for position. To achieve higher fidelity, WiDeo [36] mines backscatter reflections to realize high-resolution motion tracing.
@@ V-B2 View-Invariant Feature Extraction
Observed signal patterns often change drastically with user location and device placement. To mitigate these geometric distortions, systems like IndoTrack [41] and WiTraj [83] combine multi-link Doppler constraints to reconstruct absolute velocity vectors, correcting for view-dependent distortions. Extending this multi-link principle, RIM [13] enables RF-based inertial measurement, while WiDance [56] utilizes orthogonal links for interactive motion direction inference. Pushing further towards true domain independence, Widar3.0 [102] fuses radial velocities to derive the Body-coordinate Velocity Profile (BVP) (detailed in Sec. II-C2), enabling location-agnostic recognition. Similarly, DeepMV [92] employs a multi-view fusion mechanism to derive robust representations that are invariant to environmental variations.

### v2

@@ I-A The ISAC Paradigm and the Wi-Fi Potential
Among the various technologies supporting ISAC, commodity Wi-Fi stands out as a practical and promising platform. This prominence is primarily driven by its ubiquity and low deployment cost—intrinsic strengths that render it an ideal foundation for scalable sensing without requiring dedicated infrastructure [59, 57]. Since Channel State Information (CSI) [22, 21] became explicitly accessible, Wi-Fi has evolved into a versatile sensor supporting diverse device-free applications, ranging from localization [45, 25, 104, 84] and activity recognition [120, 75, 10, 60] to vital sign monitoring [58, 116]. Furthermore, distinct from vision-based modalities, Wi-Fi sensing inherently preserves user privacy and ensures robustness against poor lighting and severe occlusion [72, 107]. These properties establish Wi-Fi as the most accessible entry point for real-world ISAC deployments.
@@ I-D Existing Survey and Tutorial Landscape
He et al. (2020) [26] Application:
Sensing Tasks
Frames Wi-Fi sensing as “Wi-Fi vision”; categorizes perception-level tasks such as imaging, detection, and pose. Uses vision-based analogies; lacks generalizable signal-layer principles; diversity implicit but not abstracted.
Tan et al. (2022) [78] Application:
Sensing Tasks
Reviews a decade of commodity Wi-Fi sensing; taxonomizes systems by functional domains (activity, object, localization). A retrospective on application evolution; lacks bottom-up analysis of how physical-layer upgrades drive these capabilities.
Xiao et al. (2023) [101] Application:
Scenarios
Proposes a dual-axis taxonomy (task type × motion granularity); reviews diverse human sensing scenarios. High-level abstraction of application logic; lacks analysis of signal-layer capabilities or diversity mechanisms.
Chen et al. (2023) [4] Algorithm:
Frameworks
Focuses on cross-domain generalization techniques (e.g., transfer learning) for CSI-based sensing. Solves domain shift via algorithmic adaptation; overlooks how physical diversity can intrinsically reduce environmental dependency.
Yang et al. (2023) [111] Algorithm:
Frameworks
Builds SenseFi, a benchmarking library for comparing DL models across sensing tasks and datasets. Empirical emphasis on model/dataset comparison; abstracted from signal-level or physical-layer sensing implications.
Ahmad et al. (2024) [1] Algorithm:
Learning Models
Surveys Transformers and CNNs for Wi-Fi sensing; discusses training strategies and deployment challenges. Strong DL focus; limited discussion on CSI formation mechanism or physical modeling.
Radwan et al. (2025) [67] Algorithm:
Learning Models
Surveys Self-Supervised Learning (SSL) paradigms to address data scarcity; discusses pretext tasks and contrastive learning. Addresses label efficiency through learning paradigms; does not explore physical-layer enhancements to reduce data requirements.
Ma et al. (2020) [59] Signal:
Signal Processing
Surveys CSI signal modeling, preprocessing, and feature extraction pipelines. Covers signal chain depth but lacks unified diversity abstraction.
Hernandez & Bulut (2023) [32] Signal:
Signal Processing
Surveys signal processing for resource-constrained edge devices; evaluates computational efficiency. Focuses on implementation bottlenecks (computation); does not explore active physical-layer diversity.
Du et al. (2024) [13] Signal:
Standards
Comprehensive overview of the IEEE 802.11bf standard; details PHY/MAC specifications and protocols. Focuses on standard compliance and protocol definitions; does not analyze the physics regarding diversity exploitation.
@@
I-D1 Application-Centric—Sensing Tasks and Deployment Scenarios
@@
These surveys primarily organize the literature by sensing tasks or deployment scenarios, treating the physical layer as an abstracted interface. He et al. (2020) [26] frame sensing through the lens of “Wi-Fi vision,” categorizing tasks such as imaging and pose estimation to highlight RF perception capabilities. Tan et al. (2022) [78] review a decade of commodity Wi-Fi sensing, classifying systems into three functional domains: activity recognition, object sensing, and localization. Xiao et al. (2023) [101] propose a dual-axis taxonomy intersecting sensing task types with motion granularity, mapping human behaviors to system requirements. While valuable for high-level classification, these works offer limited insight into the underlying signal structure, abstracting away how physical signal diversity fundamentally constrains or enables these capabilities.
I-D2 Algorithm-Centric—Learning Models and Frameworks
These surveys focus on the computational backend, reviewing learning models while often treating Channel State Information (CSI) as a given input. Chen et al. (2023) [4] address sensing robustness, surveying cross-domain generalization techniques like transfer learning. Yang et al. (2023) [111] introduce SenseFi, a comprehensive benchmark library for evaluating deep learning models across diverse datasets. Ahmad et al. (2024) [1] provide a unified overview of modern architectures, including Transformers and CNNs, specifically for human sensing. Most recently, Radwan et al. (2025) [67] survey Self-Supervised Learning (SSL) paradigms, discussing how pretext tasks can mitigate data scarcity. However, these works generally decouple learning from signal formation. They tend to solve sensing challenges solely through model complexity, overlooking how physical-layer diversity dictates the intrinsic information content of the input data.
I-D3 Signal-Level—Signal Processing and Standards
A third line of surveys engages directly with physical layer implementation, spanning signal modeling to standardization. Ma et al. (2020) [59] systematically review the foundational signal processing pipeline, covering CSI channel modeling, noise reduction, and feature extraction. Hernandez and Bulut (2023) [32] shift the focus to edge implementation, surveying low-complexity processing techniques tailored for resource-constrained IoT devices. Du et al. (2024) [13] provide a comprehensive overview of the IEEE 802.11bf amendment, detailing the official PHY/MAC specifications for native WLAN sensing. While these works offer a holistic view of the physical layer, they predominately view the physical layer as a static set of constraints or specifications, lacking a unified abstraction on how to actively exploit physical-layer diversity to break performance limits.
@@ VI Directionality The Directional Gap:
  • Sec. VII and Sec. VIII discuss future directions and conclude the tutorial.
@@ II-A1 Physical Principles and Mathematical Representation of CSI
Channel State Information (CSI) provides a fine-grained characterization of the physical layer of a wireless channel. Conceptually, it is a discrete, sampled measurement of the underlying Channel Frequency Response (CFR), which is the continuous function that describes how a signal’s amplitude and phase change across all frequencies. In modern Wi-Fi systems based on Orthogonal Frequency-Division Multiplexing (OFDM), CSI is estimated for each subcarrier from known training sequences, such as the Long Training Field (LTF), embedded in each data packet’s preamble.
@@ II-B1 The Reality of CSI: Hardware-Induced Offsets
Modeling frameworks typically decompose these hardware-induced offsets into four major categories, each originating from specific components of the RF and baseband processing chain [103, 68]. These include Carrier Frequency Offset (CFO), Carrier Phase Offset (CPO), Sampling Frequency Offset (SFO), and Packet Detection Delay (PDD). While all four manifest as phase distortions, they differ fundamentally in their temporal characteristics and impact on sensing.
@@ II-E2 Physical Essence (Mapping Feedback to Spatial Beams)
However, stability comes at the cost of fidelity. Spatially, BFI discards amplitude information ($\bm{\Sigma}$) and absolute phase, rendering it ill-suited for attenuation-based ranging [42]. Temporally, unlike continuous CSI streams, real-world BFI is inherently sparse and bursty. It is generated only during traffic-dictated sounding intervals, leading to irregular sampling rates that challenge continuous tracking—a bottleneck addressed by advanced recovery algorithms in Sec. V-D.
@@ III-C2 Performance Evaluation and Capabilities
Figure 8: Ranging error comparison between ISAC-Fi (Monostatic) and traditional baselines (MUSIC/IFFT) [8].
@@ III-D Summary: Towards Synchronized ISAC
This chapter demonstrated how Temporal Diversity resolves the synchronization-interference dilemma. While bistatic setups suffer from clock asynchrony (losing absolute accuracy), monostatic setups face self-interference (losing visibility). To address these challenges, we categorized the engineering solutions into two paradigms:
  • Active Self-Interference Cancellation: Employs a multi-stage pipeline (analog and digital) to actively subtract transmission leakage. By recovering the complete channel response, it enables true absolute ranging for both static and dynamic targets (e.g., ISAC-Fi [8]), albeit with higher hardware complexity.
  • Physical Signal Separation: Prevents interference at the source via hardware isolation and self-mixing. It achieves high-precision Phase Coherence by eliminating clock errors, enabling robust motion tracking (e.g., SiWiS [73]). However, its reliance on DC removal generally restricts it to dynamic sensing.
Table IV compares these approaches. Ultimately, these technologies transform the Wi-Fi physical layer into a precision timing instrument, achieving Synchronized ISAC. With accurate ranging and motion tracking established, the challenge shifts to improving measurement resolution, a task governed by Frequency Diversity in the following chapter.
@@ IV-D2 The Reconstruction Challenge
Sensing Output PDP or ToA ToF of the Direct Path ToF-AoA Spectrum
@@ IV-E Summary: Towards High-Resolution ISAC
This chapter examined exploiting Frequency Diversity to realize High-Resolution ISAC. As summarized in Table V, the quest to synthesize wide effective bandwidth ($\Delta\tau=1/B$) has evolved from hardware constraints to algorithmic innovation:
  • Continuous Channel Stitching: Aggregates contiguous channels but is severely constrained by spectrum scarcity and boundary effects [103].
  • Multi-Band Frequency Sweeping: Resolves ambiguity via sequential sweeping, yet algorithmic design constraints limit its general applicability [83].
  • Discrete Channel Sampling: Synthesizes a 4.7 GHz aperture via randomized hopping and deep learning, unlocking centimeter-level resolution on commodity hardware [52].
With resolution addressed, the focus shifts to spatial coverage. The next chapter explores Link Diversity to overcome single-viewpoint occlusion through multi-perspective observations.
@@ V-B1 Omnidirectional Coverage and Geometric Triangulation
The primary objective is to eliminate inherent blind spots by aggregating views. This philosophy originates from  [114], which utilized distributed links to probabilistically infer user presence. Building on this coverage capability, WiSee [63] leveraged Doppler shifts for whole-home gesture recognition, while E-eyes [94] used fine-grained WiFi signatures from home appliances to identify activities. Beyond detection, researchers aimed for finer geometric solving. WiDraw [76] enables hands-free drawing by exploiting the AoA and hand occlusions to track fine-grained trajectories. In parallel, LiFS [90] targets room-scale localization with low human effort, modeling CSI measurements as power fading equations to solve for position. To achieve higher fidelity, WiDeo [44] mines backscatter reflections to realize high-resolution motion tracing. Most recently, Choi [11] advances this by fusing multi-link beacon CSI with inertial traces to enable unsupervised positioning without manual site surveys.
@@ V-B2 View-Invariant Feature Extraction
Observed signal patterns often change drastically with user location and device placement. To mitigate these geometric distortions, early works like CARM [92] pioneered the use of multi-link likelihood fusion to enhance recognition robustness against environmental variations. Then, systems like IndoTrack [48] and WiTraj [100] combine multi-link Doppler constraints to reconstruct absolute velocity vectors, correcting for view-dependent distortions. Extending this multi-link principle, RIM [15] enables RF-based inertial measurement, while WiDance [66] utilizes orthogonal links for interactive motion direction inference. Pushing further towards true domain independence, Widar3.0 [121] fuses radial velocities to derive the Body-coordinate Velocity Profile (BVP) (detailed in Sec. II-C2), enabling location-agnostic recognition. Similarly, DeepMV [109] employs a multi-view fusion mechanism to derive robust representations that are invariant to environmental variations.